OpenStudy (julialyxn):

Let f(x) = x − 2 and g(x) = x2 − 7x − 9. Find f(g(−1)). HELP ME

OpenStudy (phi):

g(x) is the "short-name" for a formula that uses "x" g(-1) is how we say: find the formula for g and everywhere you see x, replace the x with (-1). can you do that ?

OpenStudy (julialyxn):

so f(-1)=x-2 and g(-1) =x2-7x9?

OpenStudy (princeharryyy):